A and B working together can finish a certain job in 20/3 hours. They start the work together but after 3 hours, A became ill and B continuing the work alone finished it in 33/4 hours. How long would it take each working alone to do the job.
Answer by stanbon(75887)

A and B working together can finish a certain job in 20/3 hours. They start the work together but after 3 hours, A became ill and B continuing the work alone finished it in 33/4 hours. How long would it take each working alone to do the job.
---
Let a and b be their respective rates:
Equations:
a + b = (3/20) job/hr
3a + (3+(33/4))b = 1 job/hr
---------------------------------
Rearrange:
3a + 3b = (9/20)
3a + (45/4)b = 1
---
Subtract and solve for "b":
(33/4)b = 11/20
b = (11/20)(4/33) = (1/15)
B would take 15 hours to do the job alone
------
Solve for "a" using a + b = 3/20
a + 1/15 = 3/20
--
a = (3/20)-(1/15)
a = (9-4)/60 = 1/12
---
A would take 12 hours to do the job alone
============================================
Cheers,
Stan H.